# Artificial neural networks for magnetoencephalography: a review of an emerging field

### v3

@@ Abstract
Objective: Magnetoencephalography (MEG) is a cutting-edge neuroimaging technique that measures the intricate brain dynamics underlying cognitive processes with an unparalleled combination of high temporal and spatial precision. While MEG data analytics have traditionally relied on advanced signal processing and mathematical and statistical tools, the recent surge in artificial intelligence (AI) has led to the growing use of machine learning (ML) methods for MEG data classification. An emerging trend in this field is the use of artificial neural networks (ANNs) to address various MEG-related tasks. This review aims to provide a comprehensive overview of the state of the art in this area. Approach: This topical review included studies that applied ANNs to MEG data. Studies were sourced from PubMed, Google Scholar, arXiv, and bioRxiv using targeted search queries. The included studies were categorized into three groups: ’Classification’, ’Modeling’, and ’Other’. Key findings and trends were summarized to provide a comprehensive assessment of the field.Main results: The review identified 119 relevant studies, with 70 focused on ’Classification’, 16 on ’Modeling’, and 33 in the ’Other’ category. ’Classification’ studies addressed tasks such as brain decoding, clinical diagnostics, and BCI implementations, often achieving high predictive accuracy. ’Modeling’ studies explored the alignment between ANN activations and brain processes, offering insights into the neural representations captured by these networks. The ’Other’ category demonstrated innovative uses of ANNs for artifact correction, preprocessing, and neural source localization. Significance: By establishing a detailed portrait of the current state of the field, this review highlights the strengths and current limitations of ANNs in MEG research. It also provides practical recommendations for future work, offering a helpful reference for seasoned researchers and newcomers interested in using ANNs to explore the complex dynamics of the human brain with MEG.
@@ 1 Introduction
Artificial neural networks (ANNs) have had an enormous impact on most research fields, especially in computer vision and natural language processing ([1]). In recent years, ANNs have become increasingly used for brain data analyses and modeling, yielding ample evidence for their added value in neuroscience research ([2, 3]). Deep learning (DL) has been applied in the analysis of various types of brain imaging modalities, including structural magnetic iesonance imaging (MRI), functional magnetic resonance imaging (fMRI), electroencephalography (EEG), and magnetoencephalography (MEG). While using ANNs specifically with EEG has been the subject of comprehensive reviews (e.g. [4, 5]), using ANNs with MEG data is less common, and the current state-of-the-art has yet to be surveyed. This combination, however, is of growing interest. MEG provides direct, time-resolved recordings of brain activity, which can be used to compare the dynamics and internal representations of ANN models to those observed in the brain. For ANN researchers, this opens up possibilities for testing model outputs against high-temporal-resolution biological signals and for training models that simulate aspects of cognitive processing. For MEG researchers, artificial neural networks (ANNs) offer powerful tools to classify different brain states, uncover hidden patterns in brain activity, and model the complex structure of MEG datasets. These shared opportunities have led to an increasing number of studies exploring how MEG and ANNs can be integrated in practice. The present review seeks to fill in this gap by providing a detailed survey of the main applications of ANNs in MEG research to date. The strengths and limitations of the approaches reviewed allow us to highlight the potential of ANNs in future MEG research and the main challenges in this field.
@@ 2.1 Litterature research
2020 Lopopolo et van den Bosch [101] BCI 15 1377 CNN (4) houldout test set
@@
2021 Zhang Chuncheng et al. [107] Event 10 N/M EEGNet (4) cross-validation
@@ 3.2.1 Study aims and subcategories
Event detection studies: Lastly, this subcategory encompasses research using ANNs to identify specific, often transient, events within continuous or segmented MEG recordings. A dominant application within the reviewed literature involves the automatic detection and sometimes localization of pathological neural events, particularly epileptic spikes or high-frequency oscillations, which are crucial biomarkers for epilepsy diagnosis and treatment planning ([105, 165, 114, 116, 104, 111, 106, 113, 110, 108, 112, 109, 117]). Other studies target different types of events, such as specific visual targets within rapid presentations ([107]). Collectively, event detection studies contributed significantly to the ’Classification’ category, with epilepsy-related applications being particularly numerous (15 studies across Clinical and Event Detection subcategories combined).
@@ 3.2.2 The pipeline
Despite this shared pipeline, classification goals, model inputs, and preprocessing strategies differ substantially depending on the application, and no standard protocol has emerged across the 70 studies. Figure 4 (left panel) illustrates a schematic representation of this typical Classification pipeline, alongside comparable diagrams for the Modeling and Other application categories.
Figure 4: Representative workflows for applying Artificial Neural Networks (ANNs) to MEG data across the main application categories identified in this review. (Left) ’Classification’ pipeline: Typically involves preprocessing MEG data, feeding it to an ANN classifier, and evaluating prediction accuracy. (Center) ’Modeling’ pipeline: Often compares representations (activations) extracted from ANNs processing stimuli with corresponding preprocessed MEG data, using similarity analysis (e.g., RSA, neural predictivity) and evaluating against baselines. (Right) ’Other’ pipeline: Encompasses methodological applications like preprocessing or source localization using ANNs, often evaluated by comparison with classical approaches or metrics like AUC or spatial error.
@@ 3.2.3 The data and preprocessing
Typically, across all studies in this category, the MEG data were preprocessed before being fed as input to a neural network. This preprocessing phase generally includes down-sampling, band-pass filtering, and either de-noising and/or artifact removal. Of the 70 studies in this category, 12 studies (Decoding: [57, 61, 53, 50] BCI: [91, 97, 100, 98, 92, 93], clinical: [86], event detection: [116]) use discrete wavelet transform to increase signal-to-noise ratio (SNR) in the data as a preprocessing step ([170]). Notably, these 12 studies are spread across only four distinct research groups representing a small portion of the reviewed studies. Five studies employ alternative de-noising techniques ([62, 69, 89, 90, 165]). Additionally, 16 studies use ICA to remove heart and/or eye movement artefacts ([54, 71, 67, 95, 88, 84, 105, 107, 89, 90, 74, 68, 82, 78, 83, 56]). The remaining 38 studies do not explicitly mention de-noising or artifact removal techniques in their preprocessing pipeline.
@@ 3.2.6 Imbalanced datasets
The limited attention given to this issue is concerning, given the known sensitivity of neural networks to unbalanced data, particularly when applied to high-dimensional signals like MEG. Future work in this area would benefit from more systematic reporting of class distributions, explicit justification of performance metrics, and comparative evaluations of different strategies for mitigating imbalance ([184]). Doing so would improve the interpretability and reproducibility of results and allow for more meaningful comparisons across studies.
@@ 3.4.4 Network architectures
In source localization studies, architectures that explicitly modeled temporal dynamics were more common. These included Long Short-Term Memory (LSTM) networks and other types of Recurrent Neural Networks (RNNs), which are well-suited to tracking changes in MEG activity over time. Some studies also explored ResNet-based models [137] or more novel frameworks such as the multiple penalized state space (MPSS) architecture, designed to integrate spatial and temporal priors [168]. A few studies combined recurrent and convolutional elements to capture both local and distributed features in the MEG signals.
@@ 3.4.5 Training and validation techniques
Validation strategies in these studies included simple train/validation/test splits, K-fold cross-validation, and Leave-One-Subject-Out (LOSO) schemes. LOSO was most commonly used when generalizing across participants was crucial or when datasets were small. K-fold and bootstrap resampling methods were employed to ensure robustness, particularly when evaluating architectural innovations across multiple configurations. In methods papers proposing general-purpose architectures, performance was often tested across different tasks or datasets to demonstrate transferability.
@@ 4.3.2 Model validation techniques
In many neural network applications, the vast amount of data combined with the typical training duration for ANNs makes cross-validation an impractical method for model validation. Instead, most of the time, the dataset is randomly split into three subsets to validate the model: training, validation, and test (also called evaluation set). The training set generally represents 50 to 80 percent of the initial dataset, and the remaining data is split in half to create the validation and test sets. The training set is employed to train the network with a set of hyperparameters and tested on the validation set. The test set is only used at the end when the architecture and hyper-parameters are fixed and once the desired performance has been reached on the validation set. It is solely used to evaluate the final model performance. However, when data is scarce, as it is often the case for neuroscience studies, it is recommended to use re-sampling methods such as bootstrapping or cross-validation to make use of the data as much as possible ([177]). It will allow the train and validation sets to be changed several times to acquire a less noisy measure of the model’s performance. In the studies included in this survey, when cross-validation is used, K-fold was the preferred technique, with the value of K ranging from 4 to 50. When the goal is to test generalization to new subjects, the "Leave P subjects out" cross-validation or a stratified version of K-Fold is recommended.
@@ 4.3.3 Model performance evaluation and baseline comparisons
Some classification studies did not include baseline results against which one could have compared the reported ANN performances. This can be a relevant concern given that classical shallow ML approaches (e.g. logistic regression [201], random forest [179], SVM [202], etc.) often involve simpler classifier implementations and potentially faster training times —once appropriate features have been engineered—, compared to designing and training complex ANNs. However, the overall practical advantage depends heavily on the computational complexity of the necessary feature extraction pipeline, which can itself be intensive and time-consuming, potentially offsetting the gains in classifier simplicity. Additionally, it is critical to evaluate classification performance against chance-level baselines, particularly in studies with small sample sizes, where theoretical chance levels (e.g., 50% accuracy for two-class problems) can be misleadingly optimistic. Limited data can lead to inflated accuracies, even exceeding 70%, purely by chance and rigorous statistical evaluations, such as permutation tests, are required to properly establish meaningful baseline performances, thus ensuring robust assessments of classifier significance [203]. Generally, one should always evaluate simple solutions before spending time and effort on more complex ones, which might also be more challenging to interpret. Similarly, only a few articles among the modeling studies mentioned some kind of baseline performance comparison ([121, 122, 118, 119, 126, 129]). Ideally, modeling studies should report NC and/or evaluate the performance of their similarity analysis with an untrained model. Together, these measures would reinforce robust similarity scores’ reliability and help identify and discard weaker results.
@@ 4.3.5 Variability in MEG data acquisition, preprocessing, and experimental protocols
One of the main challenges in applying artificial neural networks (ANNs) to MEG research is the variability in acquisition hardware, preprocessing steps, and experimental protocols across studies. These differences can affect data quality and consistency, which in turn influence model performance and the ability to compare results across studies.
@@
To address these challenges, future studies would benefit from more consistent preprocessing pipelines, standardized task protocols where feasible, and clearer reporting of methods. Efforts like BIDS-MEG and open datasets with thorough documentation can support better reproducibility and enable more reliable benchmarking across studies.
@@ 4.3.6 The issue of reproducibility
The utility of these public datasets for reproducible research and cross-study comparisons is greatly enhanced by standardized data formats and organizational structures, notably the Brain Imaging Data Structure extension for MEG (MEG-BIDS) [208]. Resources providing standardized, open-access MEG recordings suitable for ANN analysis include repositories and datasets such as OpenNeuro (https://openneuro.org/), the human connectome project (https://www.humanconnectome.org/), the Cambridge centre for ageing and neuroscience (Cam-CAN) dataset (https://camcan-archive.mrc-cbu.cam.ac.uk/dataaccess/), the open MEG archive (OMEGA) database (https://www.mcgill.ca/bic/neuroinformatics/omega), and the MNE sample datasets (https://mne.tools/stable/overview/datasets_index.html), offering valuable benchmarks for training, testing, and comparing models.
@@ 4.3.7 Interpretation and visualizations
In BCI studies, where the primary concern is the algorithm’s performance, the black-box nature of ANNs is often considered less problematic. However, in many other applications, including brain decoding studies, gaining insight into what is happening in the model’s latent space is key to understanding how the information was extracted from the signal, as well as why and how decoding is possible. In domains like image classification or object detection, there are various tools (Grad-CAM [209], tSNE [210], UMAP [211]) that allow the user to have a representation of what the network ’sees’ and what properties of the data are important in its decision-making process. A large proportion (52 out of 70) of the classification studies, where interpretation would be the most interesting, did not use these tools to gain insights into the representations learned and used by the model. A similar pattern of limited application was observed in the ’Other’ category studies reviewed. While interpretability might be considered less critical for some preprocessing or source localization tasks compared to decoding, it remains vital for evaluating novel methodological approaches. Indeed, roughly half of the ’Methods’ studies (9 out of 16) included visualization techniques such as filter or activation map analysis. Interpretability tools were also used in a majority of the ’Preprocessing’ studies (5 out of 8), often employing methods like Grad-CAM [182]. However, interpretability techniques were notably absent in all reviewed ANN-based source localization studies. This suggests that developing and applying robust interpretability methods beyond classification and basic modeling tasks remains an important area for future work in MEG research using ANNs.
@@ Acknowledgements
I.R. received support from the Canada CIFAR AI chair program and the Canada excellence research chairs program. K.J. was supported by Canada research chairs program funding (950-232368). V.H. was supported through the Canadian Institutes of health tesearch project grant (166197). H.A. was funded through the faculty of medicine at the university of Montreal’s merit scholarship. A.D. was funded through computational neuroscience and cognitive neuroimaging lab (CoCoLab) and Montreal institute of learning algorithms (MILA) Scholarships.

### v4

@@ Abstract
Objective: Magnetoencephalography (MEG) is a cutting-edge neuroimaging technique that measures the intricate brain dynamics underlying cognitive processes with an unparalleled combination of high temporal and spatial precision. While MEG data analytics have traditionally relied on advanced signal processing and mathematical and statistical tools, the recent surge in artificial intelligence (AI) has led to the growing use of machine learning (ML) methods for MEG data classification. An emerging trend in this field is the use of artificial neural networks (ANNs) to address various MEG-related tasks. This review aims to provide a comprehensive overview of the state of the art in this area. Approach: This topical review included studies that applied ANNs to MEG data. Studies were sourced from PubMed, Google Scholar, arXiv, and bioRxiv using targeted search queries. The included studies were categorized into three groups: ’Classification’, ’Modeling’, and ’Other’. Key findings and trends were summarized to provide a comprehensive assessment of the field.Main results: We identified 119 relevant studies, with 70 focused on ’Classification’, 16 on ’Modeling’, and 33 in the ’Other’ category. ’Classification’ studies addressed tasks such as brain decoding, clinical diagnostics, and BCI implementations, often achieving high predictive accuracy. ’Modeling’ studies explored the alignment between ANN activations and brain processes, offering insights into the neural representations captured by these networks. The ’Other’ category demonstrated innovative uses of ANNs for artifact correction, preprocessing, and neural source localization. Significance: By establishing a detailed portrait of the current state of the field, this review highlights the strengths and current limitations of ANNs in MEG research. It also provides practical recommendations for future work, offering a helpful reference for seasoned researchers and newcomers interested in using ANNs to explore the complex dynamics of the human brain with MEG.
@@ 1 Introduction
Artificial neural networks (ANNs) have had an enormous impact on most research fields, especially in computer vision and natural language processing ([1]). In recent years, ANNs have become increasingly used for brain data analyses and modeling, yielding ample evidence for their added value in neuroscience research ([2, 3]). Deep learning (DL) has been applied in the analysis of various types of brain imaging modalities, including structural magnetic iesonance imaging (MRI), functional magnetic resonance imaging (fMRI), electroencephalography (EEG), and magnetoencephalography (MEG). While using ANNs specifically with EEG has been the subject of comprehensive reviews (e.g. [4, 5]), using ANNs with MEG data is less common, and the current state-of-the-art has yet to be surveyed. This combination, however, is of growing interest. MEG provides direct, time-resolved recordings of brain activity, which can be used to compare the dynamics and internal representations of ANN models to those observed in the brain. For ANN researchers, this opens up possibilities for testing model outputs against high-temporal-resolution biological signals and for training models that simulate aspects of cognitive processing. For MEG researchers, ANNs offer powerful tools to classify different brain states, uncover hidden patterns in brain activity, and model the complex structure of MEG datasets. These shared opportunities have led to an increasing number of studies exploring how MEG and ANNs can be integrated in practice. The present review seeks to fill in this gap by providing a detailed survey of the main applications of ANNs in MEG research to date. The strengths and limitations of the approaches reviewed allow us to highlight the potential of ANNs in future MEG research and the main challenges in this field.
@@ 2.1 Litterature research
2020 Lopopolo and van den Bosch [101] BCI 15 1377 CNN (4) houldout test set
@@
2021 Zhang et al. [107] Event 10 N/M EEGNet (4) cross-validation
@@ 3.2.1 Study aims and subcategories
Event detection studies: Lastly, this subcategory encompasses research using ANNs to identify specific, often transient, events within continuous or segmented MEG recordings. A dominant application within the reviewed literature involves the automatic detection and sometimes localization of pathological neural events, particularly epileptic spikes or high-frequency oscillations, which are crucial biomarkers for epilepsy diagnosis and treatment planning ([105, 113, 115, 117, 104, 111, 106, 114, 110, 108, 112, 109, 118]). Other studies target different types of events, such as specific visual targets within rapid presentations ([107]). Collectively, event detection studies contributed significantly to the ’Classification’ category, with epilepsy-related applications being particularly numerous (15 studies across ’Clinical’ and ’Event detection’ subcategories combined).
@@ 3.2.2 The pipeline
Despite this shared pipeline, classification goals, model inputs, and preprocessing strategies differ substantially depending on the application, and no standard protocol has emerged across the 70 studies. Figure 4 (left panel) illustrates a schematic representation of this typical ’Classification’ pipeline, alongside comparable diagrams for the ’Modeling’ and ’Other’ application categories.
Figure 4: Representative workflows for applying ANNs to MEG data across the main application categories identified in this review. (left) ’Classification’ pipeline: typically involves preprocessing MEG data, feeding it to an ANN classifier, and evaluating prediction accuracy. (center) ’Modeling’ pipeline: often compares representations (activations) extracted from ANNs processing stimuli with corresponding preprocessed MEG data, using similarity analysis (e.g., RSA, neural predictivity) and evaluating against baselines. (right) ’Other’ pipeline: encompasses methodological applications like preprocessing or source localization using ANNs, often evaluated by comparison with classical approaches or metrics like AUC or spatial error.
@@ 3.2.3 The data and preprocessing
Typically, across all studies in this category, the MEG data were preprocessed before being fed as input to a neural network. This preprocessing phase generally includes down-sampling, band-pass filtering, and either de-noising and/or artifact removal. Of the 70 studies in this category, 12 studies (’decoding’: [57, 61, 53, 50] ’BCI’: [91, 97, 100, 98, 92, 93], ’clinical’: [86], ’event detection’: [117]) use discrete wavelet transform to increase signal-to-noise ratio (SNR) in the data as a preprocessing step ([169]). Notably, these 12 studies are spread across only four distinct research groups representing a small portion of the reviewed studies. Five studies employ alternative de-noising techniques ([62, 69, 89, 90, 113]). Additionally, 16 studies use ICA to remove heart and/or eye movement artefacts ([54, 71, 67, 95, 88, 84, 105, 107, 89, 90, 74, 68, 82, 78, 83, 56]). The remaining 38 studies do not explicitly mention de-noising or artifact removal techniques in their preprocessing pipeline.
@@ 3.2.6 Imbalanced datasets
The limited attention given to this issue is concerning, given the known sensitivity of neural networks to imbalanced data, particularly when applied to high-dimensional signals like MEG. Future work in this area would benefit from more systematic reporting of class distributions, explicit justification of performance metrics, and comparative evaluations of different strategies for mitigating imbalance ([184]). Doing so would improve the interpretability and reproducibility of results and allow for more meaningful comparisons across studies.
@@ 3.4.4 Network architectures
In source localization studies, architectures that explicitly modeled temporal dynamics were more common. These included long short-term memory (LSTM) networks and other types of RNNs, which are well-suited to tracking changes in MEG activity over time. Some studies also explored ResNet-based models [138] or more novel frameworks such as the multiple penalized state space (MPSS) architecture, designed to integrate spatial and temporal priors [167]. A few studies combined recurrent and convolutional elements to capture both local and distributed features in the MEG signals.
@@ 3.4.5 Training and validation techniques
Validation strategies in these studies included simple train/validation/test splits, K-fold cross-validation, and LOSO schemes. LOSO was most commonly used when generalizing across participants was crucial or when datasets were small. K-fold and bootstrap resampling methods were employed to ensure robustness, particularly when evaluating architectural innovations across multiple configurations. In method papers proposing general-purpose architectures, performance was often tested across different tasks or datasets to demonstrate transferability.
@@ 4.3.2 Model validation techniques
In many neural network applications, the vast amount of data combined with the typical training duration for ANNs makes cross-validation an impractical method for model validation. Instead, most of the time, the dataset is randomly split into three subsets to validate the model: training, validation, and test (also called evaluation set). The training set generally represents 50 to 80 percent of the initial dataset, and the remaining data is split in half to create the validation and test sets. The training set is employed to train the network with a set of hyperparameters and tested on the validation set. The test set is only used at the end when the architecture and hyper-parameters are fixed and once the desired performance has been reached on the validation set. It is solely used to evaluate the final model performance. However, when data is scarce, as it is often the case for neuroscience studies, it is recommended to use re-sampling methods such as bootstrapping or cross-validation to make use of the data as much as possible ([177]). It will allow the train and validation sets to be changed several times to acquire a less noisy measure of the model’s performance. In the studies included in this survey, when cross-validation is used, K-fold was the preferred technique, with the value of K ranging from 4 to 50. When the goal is to test generalization to new subjects, the leave P subjects out (LOPO) cross-validation or a stratified version of K-Fold is recommended.
@@ 4.3.3 Model performance evaluation and baseline comparisons
Some classification studies did not include baseline results against which one could have compared the reported ANN performances. This can be a relevant concern given that classical shallow ML approaches (e.g. logistic regression [201], random forest [179], SVM [202], etc.) often involve simpler classifier implementations and potentially faster training times —once appropriate features have been engineered—, compared to designing and training complex ANNs. However, the overall practical advantage depends heavily on the computational complexity of the necessary feature extraction pipeline, which can itself be intensive and time-consuming, potentially offsetting the gains in classifier simplicity. Additionally, small sample size can lead to inflated accuracies, that can substantially exceed the theoretical chance level of 50% for two classes (eg. reaching 70% purely by chance). Therefore, rigorous statistical evaluations, such as permutation tests are required to properly establish meaningful baseline performances, thus ensuring robust assessments of classifier significance [203]. Generally, one should always evaluate simple solutions before spending time and effort on more complex ones, which might also be more challenging to interpret. However, only a few articles among the modeling studies mentioned some kind of baseline performance comparison ([122, 123, 119, 120, 127, 130]). Ideally, modeling studies should report NC and/or evaluate the performance of their similarity analysis with an untrained model. Together, these measures would reinforce robust similarity scores’ reliability and help identify and discard weaker results.
@@ 4.3.5 Variability in MEG data acquisition, preprocessing, and experimental protocols
One of the main challenges in applying ANNs to MEG research is the variability in acquisition hardware, preprocessing steps, and experimental protocols across studies. These differences can affect data quality and consistency, which in turn influence model performance and the ability to compare results across studies.
@@
To address these challenges, future studies would benefit from more consistent preprocessing pipelines, standardized task protocols where feasible, and clearer reporting of methods. Efforts such as the brain imaging data structure extension for MEG (MEG-BIDS [205]) and open datasets with thorough documentation can support better reproducibility and enable more reliable benchmarking across studies.
@@ 4.3.6 The issue of reproducibility
The utility of these public datasets for reproducible research and cross-study comparisons is greatly enhanced by standardized data formats and organizational structures, such as MEG-BIDS ([205]). Resources providing standardized, open-access MEG recordings suitable for ANN analysis include repositories and datasets such as OpenNeuro (https://openneuro.org/), the human connectome project (https://www.humanconnectome.org/), the Cambridge centre for ageing and neuroscience (Cam-CAN) dataset (https://camcan-archive.mrc-cbu.cam.ac.uk/dataaccess/), the open MEG archive (OMEGA) database (https://www.mcgill.ca/bic/neuroinformatics/omega), and the MNE sample datasets (https://mne.tools/stable/overview/datasets_index.html), offering valuable benchmarks for training, testing, and comparing models.
@@ 4.3.7 Interpretation and visualizations
In BCI studies, where the primary concern is the algorithm’s performance, the black-box nature of ANNs is often considered less problematic. However, in many other applications, including brain decoding studies, gaining insight into what is happening in the model’s latent space is key to understanding how the information was extracted from the signal, as well as why and how decoding is possible. In domains like image classification or object detection, there are various tools (Grad-CAM [209], tSNE [210], UMAP [211]) that allow the user to have a representation of what the network ’sees’ and what properties of the data are important in its decision-making process.
A large proportion (52 out of 70) of the classification studies, where interpretation would be the most interesting, did not use these tools to gain insights into the representations learned and used by the model. A similar pattern of limited application was observed in the ’Other’ category studies reviewed. While interpretability might be considered less critical for some preprocessing or source localization tasks compared to decoding, it remains vital for evaluating novel methodological approaches. Indeed, roughly half of the ’Methods’ studies (9 out of 16) included visualization techniques such as filter or activation map analysis. Interpretability tools were also used in a majority of the ’Preprocessing’ studies (5 out of 8), often employing methods like Grad-CAM [182]. However, interpretability techniques were notably absent in all reviewed ANN-based source localization studies. This suggests that developing and applying robust interpretability methods beyond classification and basic modeling tasks remains an important area for future work in MEG research using ANNs.
@@ Acknowledgements
I.R. received support from the Canada CIFAR AI chair program and the Canada excellence research chairs program. K.J. was supported by Canada research chairs program funding (950-232368). V.H. was supported through the ianadian institutes of health tesearch project grant (166197). H.A. was funded through the faculty of medicine at the university of Montreal’s merit scholarship. A.D. was funded through computational neuroscience and cognitive neuroimaging lab (CoCoLab) and Montreal institute of learning algorithms (MILA) scholarships.